OpenStudy (anonymous):

Please help!!!!! Will fan and Medal :) Choose the answer. 4. (3 pt) Which equation demonstrates the associative property? A. –(33 + 4) = –33 + (–4) B. (–16 + 9) + (–19) = –16 + (9 + (–19)) C. (–5 + 11) + 3 = (11 + (–5)) + 3 D. 2 + (–24) = –24 + 2

OpenStudy (anonymous):

associative law: (x * y) * z = x * (y * z) for any x, y, z * is any mathematical operation

OpenStudy (anonymous):

so whats ur answer? u tell.. will confirm u

OpenStudy (anonymous):

to be honest i'm still a bit confused :/

OpenStudy (anonymous):

okk in simple language... only brackets " ( ) " change : that is : (a+b)+c=a+(b+c)

OpenStudy (anonymous):

thank you, hold on i'll work it out and be right back :)

OpenStudy (anonymous):

@vedic A?

OpenStudy (anonymous):

no its (B) dear... sorry had meeting...